OpenStudy (anonymous):

What is the surface area of this rectangular prism? A. 209 in2 B. 836 in2 C. 418 in2 D. 520 in2

OpenStudy (anonymous):

A = 2 (area of base) + (perimeter of base) * Height

OpenStudy (anonymous):

 im not good at drawing >.<

OpenStudy (anonymous):

d?

OpenStudy (nurali):

Area = 2ab + 2bc + 2ac

OpenStudy (anonymous):

id get it..

OpenStudy (timaashorty):

\[= 2(8*13) + 2(8*5)+ 2(13*5)\] \[= 2(104)+ 2(45)+2(65)\] \[= 208 + 90 + 130\] \[Area= 418 in ^{2}\]

OpenStudy (anonymous):

ty so vm

OpenStudy (timaashorty):

your welcome